OpenStudy (anonymous):

Kristin opens a savings account with $3500. She deposits $2400 every year into the account that has a 0.65% interest rate, compounded quarterly. If she doesn't withdraw any money, what will the account balance be in 10 years?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

$28,551.51 $28,511.25 $24,816.65 $24,776.39

jimthompson5910 (jim_thompson5910):

this is a future value of annuity problem do you have a formula for that?

OpenStudy (anonymous):

yes i think so. let me find it real quick

OpenStudy (anonymous):

is this it?

jimthompson5910 (jim_thompson5910):

yes, FVOA = future value of ordinary annuity

jimthompson5910 (jim_thompson5910):

I just need to figure out what to do with the initial $3500

OpenStudy (anonymous):

thats what confused me

jimthompson5910 (jim_thompson5910):

yeah it's throwing me off too, but I'm looking for ways to address it. One moment

OpenStudy (anonymous):

take your time, thank you for trying

OpenStudy (anonymous):

could it be a combination of lump sum and ordinary annuity?

OpenStudy (anonymous):

probably not but i'm just thinking of different equations

jimthompson5910 (jim_thompson5910):

that's possible you may have to use FV = PV*(1+r/n)^(n*t) to figure out how much money Kristin would have if she did NOT deposit any money at all. This is just for the $3500 that's already in there. and then you'd add on the annuity part

OpenStudy (anonymous):

alright i'll see what i get

OpenStudy (anonymous):

i got 3734.86 would i solve the FVOA equation and just add them together?

OpenStudy (anonymous):

nope it came out way too high

jimthompson5910 (jim_thompson5910):

well we at least know the answer is NOT around 24,000 because over 10 years, Kristin would deposit 10*2400 = 24000 dollars on top of the 3500 already there so it has to be either A or B

OpenStudy (anonymous):

so should i just guess?

jimthompson5910 (jim_thompson5910):

no, let me still try, one sec

OpenStudy (anonymous):

oh alright thank you

OpenStudy (anonymous):

this class is kicking my butt :(

jimthompson5910 (jim_thompson5910):

do they offer any other future value formula?

OpenStudy (anonymous):

only for lump sum ---> FV = PV(1 + i)nt

jimthompson5910 (jim_thompson5910):

I'm looking through here http://faculty.atu.edu/mfinan/actuarieshall/Formulas.pdf to see if I can find anything

OpenStudy (anonymous):

she that looks like Spanish to me haha i didn't understand any of it.

jimthompson5910 (jim_thompson5910):

I think I figured it out, but I'll need a bit of time

OpenStudy (anonymous):

take your time

jimthompson5910 (jim_thompson5910):

sorry this is taking forever, but what I'm doing now is trying out different values and methods in excel

jimthompson5910 (jim_thompson5910):

I'm trying to get to either $28,551.51 or $28,511.25